# Yuhong Liu, Yong Yang, Di Wang, Jie Jiang, Chunchao Guo

### v2

@@ 5 Results
Adaptive CFG. We verify the effectiveness of adaptive classifier-free guidance (CFG) on generated multi-view images in Fig. 6. Traditional fixed CFG throughout the denoising process often results in dark shadows in the back views. While the time-adaptive CFG introduced by Consistent123 [56] helps mitigate the shadow issue, it overlooks the relationships between views. In our camera orbit settings, the condition image has more area visible from the front view. The slow CFG would reduce the condition control for the front view, while the high CFG leaves excessive control over the back view, causing the model to replicate details from the front, such as the logo on the back of the cup. By dynamically adjusting the CFG during the generation process, we achieve a significant improvement in image quality. The adaptive CFG mechanism effectively prevents oversaturation and enables the model to generate more coherent and realistic multi-view images.
We evaluate the effectiveness of adaptive classifier-free guidance (CFG) on the generated multi-view images, as shown in Fig. 6. Traditional fixed CFG throughout the denoising process often tends to generate dark shadows in the back views. Although the time-adaptive CFG introduced by Consistent123 [56] helps mitigate this shadowing issue, it ignores the relationships between views. In our camera orbit settings, the condition image has a larger visible area from the front view. A low CFG reduces the control for the front view generation, while a high CFG exerts too much control over the back view generation, causing the model to replicate details from the front, such as the copied logo on the back of the cup. By dynamically adjusting the CFG during the generation process, we achieve a balance between controllability and diversity across different views, enabling the model to produce more coherent and realistic multi-view images.

### v3

@@ 5 Results
Adaptive CFG. We evaluate the effectiveness of adaptive classifier-free guidance (CFG) on the generated multi-view images, as shown in Fig. 6. Traditional fixed CFG throughout the denoising process often tends to generate dark shadows in the back views. Although the time-adaptive CFG introduced by Consistent123 [56] helps mitigate this shadowing issue, it ignores the relationships between views. In our camera orbit settings, the condition image has a larger visible area from the front view. A low CFG reduces the control for the front view generation, while a high CFG exerts too much control over the back view generation, causing the model to replicate details from the front, such as the copied logo on the back of the cup. By dynamically adjusting the CFG according to viewpoint distance during the generation process, we achieve a balance between controllability and diversity across different views, enabling the model to produce more coherent and realistic multi-view images.